OpenStudy (anonymous):

Technically Chem but it just involves calculations, so... Could someone show me how to convert them to the required units? A block of material has a volume of 2 m3. It has a mass of 1 × 104 kg. Calculate the density of the material sample and express your results in g/cm3 with appropriate scientific notation.

OpenStudy (anonymous):

\[104kg * \frac{ 1000 Grams }{ 1kg } = 104000 Grams\] \[2 m^3 * (\frac{ 100 cm }{ 1m })^3 = 2000000 cm^3\] -- Remember to cube that 100 cm! \[\frac{ 104000 Grams }{ 2000000 cm^3 } = .052 g/cm^3\] Now for appropriate scientific notation = 5.2 x 10^-2 Any problems please ask.

OpenStudy (anonymous):

Thank you so much! I really appreciate the explanation. :)

OpenStudy (anonymous):

Looking at your problem again I may have mistook something so please double check it! Is your kg supposed to be 1 * 10^4 ? I might have mistook it for simply 104, but now you know how to do it so yea!